# Stereo4D: Learning How Things Move in 3D from Internet Stereo Videos

### v1

@@ 9.1 Equirectangular videos
The raw videos that we collect (see examples in Fig. 15) are natively stored in a cropped equirectangular format, which differs from a full 360 equirectangular projection as the horizontal field of view of the cropped format typically spans 180—half of a full sphere. These videos often contain metadata specifying the horizontal and vertical field of view. For instance, metadata for a typical video might specify $\mathsf{start_{yaw}}=-90.0^{\circ}$, $\mathsf{end_{yaw}}=90.0^{\circ}$, $\mathsf{start_{tilt}}=-90.0^{\circ}$, $\mathsf{end_{tilt}}=90.0^{\circ}$; Since many VR180 videos are designed for an immersive VR experience, they are typically viewed with headsets. Hence, the baseline between the left and right cameras typically closely matches the average human eye distance of 6.3 cm.
9.2 SfM
@@ 9.3 Depth estimation
We extract long-range 2D point trajectories using BootsTAP [17]. We run tracking on the left-eye video only. For every 10 frames, we uniformly initialize query points on image with stride 4. We then remove duplicated queries if earlier tracks fall within 1 pixel of a query point.
9.5 Choice of FoV and resolution for perspective projection.
@@
10 DynaDUSt3R training details.
@@
Training. The network takes input the two RGB images as well as query times $t_{q}=\{0,1,\frac{t_{\mathsf{aux}}-t_{0}}{t_{1}-t_{0}}\}$ and predicts the pointmaps for the two input views and motionmaps for each query $t_{q}$. We supervise the network with losses defined in Eqn. 6 and 7. We initialize our network with the DUSt3R weights and initialize the motion head with the same weights as the point head. We finetune for 49k iterations with batch size 64, learning rate $2.5\times 10^{-5}$, and optimized by Adam with weight decay 0.95.

### v2

@@ 10.1 Equirectangular videos
The raw videos that we collect (see examples in Fig. 18) are natively stored in a cropped equirectangular format, which differs from a full 360 equirectangular projection as the horizontal field of view of the cropped format typically spans 180—half of a full sphere. These videos often contain metadata specifying the horizontal and vertical field of view. For instance, metadata for a typical video might specify $\mathsf{start_{yaw}}=-90.0^{\circ}$, $\mathsf{end_{yaw}}=90.0^{\circ}$, $\mathsf{start_{tilt}}=-90.0^{\circ}$, $\mathsf{end_{tilt}}=90.0^{\circ}$; Since many VR180 videos are designed for an immersive VR experience, they are typically viewed with headsets. Hence, the baseline between the left and right cameras typically closely matches the average human interpupillary distance of 6.3 cm.
10.2 Structure from motion
@@ 10.3 Depth estimation
We extract long-range 2D point trajectories using Boots-TAP [17]. We run tracking on the left-eye video only. For every 10 frames, we uniformly initialize query points on image with stride 4. We then remove duplicated queries if earlier tracks fall within 1 pixel of a query point.
10.5 Choice of FoV and resolution for perspective projection
@@
11 DynaDUSt3R training details
@@
Training. The network takes input the two RGB images as well as query times $t_{q}=\{0,1,\frac{t_{\mathsf{aux}}-t_{0}}{t_{1}-t_{0}}\}$ and predicts the pointmaps for the two input views and motionmaps for each query $t_{q}$. We supervise the network with losses defined in Eqn. 6 and 7. We initialize our network with the DUSt3R weights and initialize the motion head with the same weights as the point head. We finetune for 49k iterations with batch size 64, learning rate $2.5\times 10^{-5}$, and optimized by Adam with weight decay 0.05.